OpenStudy (anonymous):

-2x-1=3x/+5/2 solve for X

OpenStudy (mertsj):

\[-2x-1=\frac{3x}{\frac{5}{2}}\]

OpenStudy (mertsj):

Is that the problem or have you posted it incorectly?

OpenStudy (anonymous):

not it

OpenStudy (anonymous):

i am back

OpenStudy (mertsj):

Well, it's great that you are back but it is still impossible to tell what the problem is so why don't you repost it correctly?